OpenStudy (anonymous):

are the given vectors normal? a = (5, -2) and b = (6, 15)

OpenStudy (anonymous):

take a dot product . if dot product=0 then is

OpenStudy (anonymous):

*then yes

OpenStudy (anonymous):

@modphysnoob so (5 x 6) + (-2 x 15) = 30 + (-30) = 0. so they are normal, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so they are normal